a particle of mass 4kg moving with a velocity of 2 m/s collides elastically with particle of mass 6kg moving with velocity of -4m/s.what is the centre of mass?
1
Expert's answer
2020-04-09T09:31:19-0400

As per the given question,

Mass of the particle "(m_1)=4kg"

Velocity of the object "(v_1)=2 m\/sec"

Mass of the second "(m_2)=6kg"

Velocity of the second object "(v_2)=-4m\/sec"

So, "V_{com}=\dfrac{m_1v_2+m_2v_2}{m_1+m_2}"


"\Rightarrow V_{com}=\dfrac{4\times 2-6\times 4}{4+6}"


"\Rightarrow V_{com}=\dfrac{8-24}{10}=\dfrac{-16}{10}=-1.6m/sec"

Hence the velocity of center of mass will be "V_{com}=-1.6m\/sec"